OpenStudy (anonymous):

Give a formula for your yearly salary as a function of time (t) in years if you are given a base salary of $40,000 and a 2.5% raise per year

OpenStudy (phi):

any idea?

OpenStudy (anonymous):

Well, my original idea was that it would be like an interest equation and the amount added to the salary each year could be calculated by doing 40,000* 0.025*T

OpenStudy (phi):

good idea, but do you know the formula for compound interest?

OpenStudy (anonymous):

S= P (1+r/t)^t correct?

OpenStudy (phi):

I would use S= P(1+i)^t in this case, where r is a yearly rate, and things are compounded every year.

OpenStudy (phi):

so S= 40000*1.025^t where t is in years

OpenStudy (anonymous):

Thanks for the help! That is the same equation that I get using S= P (1+r/t)^t